Find the inverse of the function.

f(x) = 7x - 1
1
Expert's answer
2014-09-08T13:47:33-0400
Let y=7x-1. Express x in terms of y: y+1=7x, hence x=(y+1)/7. Thus, function g(x)=(x+1)/7 is the inverse function of f(x)=7x-1